OpenStudy (imnotarobot01):

HELP ME PLEASE (cos Θ + cos Θ)2 + (cos Θ + cos Θ)2

OpenStudy (imnotarobot01):

sin2 Θ 4 8cos2Θ cos2 Θ

OpenStudy (sadlalkdja12111312):

what your answer and use a calculator

OpenStudy (imnotarobot01):

i don't know how to do it

OpenStudy (imnotarobot01):

explain how and then i can figure it out

OpenStudy (sadlalkdja12111312):

cos Θ + cos Θ +2 + cos Θ + cos +2

OpenStudy (sachintha):

Is it like this \((cos\,\theta+cos\,\theta)^2+(cos\,\theta+cos\,\theta)^2\) ?

OpenStudy (imnotarobot01):

yes exactly

OpenStudy (imnotarobot01):

and i posted the answers above , can you help me out @Sachintha

OpenStudy (sadlalkdja12111312):

i dont know it was square of what in the bracket

OpenStudy (imnotarobot01):

you're not making sense if you don't know how to do it, don't confuse yourself. ill find someone else who knows it

OpenStudy (michele_laino):

hint: \[\left( {\cos \theta + \cos \theta } \right) = {\left( {2\cos \theta } \right)^2} = 4{\left( {\cos \theta } \right)^2}\]

OpenStudy (michele_laino):

oops.. I have made a typo: \[{\left( {\cos \theta + \cos \theta } \right)^2} = {\left( {2\cos \theta } \right)^2} = 4{\left( {\cos \theta } \right)^2}\]

jhonyy9 (jhonyy9):

hope you think 4cos^2 theta - yes ?

OpenStudy (michele_laino):

yes!

OpenStudy (imnotarobot01):

yes i was just writing that :)

jhonyy9 (jhonyy9):

just bc. in these above wrote choices are wrote in this way - sorry for ...

OpenStudy (imnotarobot01):

What answer would be correct them, B?

OpenStudy (michele_laino):

I think it is option C, since the function is: \[f\left( \theta \right) = {\left( {\cos \theta + \cos \theta } \right)^2} + {\left( {\cos \theta + \cos \theta } \right)^2}\]

OpenStudy (imnotarobot01):

Okay thank you again :)

OpenStudy (michele_laino):

:)

OpenStudy (imnotarobot01):

Can you help me with ne more